### v1

@@ 6 Methods for Continuous Dynamic HGR
In contrast to isolated dynamic hand gesture recognition (IDHGR), where each sample contains at most one gesture, continuous hand gesture recognition (CHGR) deals with unsegmented gesture sequences. The goal is to recognize gestures from untrimmed videos and generate the corresponding gesture sequence. CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss [188, 55].
While CHGR has been explored in domains like autonomous driving [45, 185], most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss. CSLR aims to produce a sequence of glosses [^2] Glosses are atomic lexical units in sign language, analogous to words in spoken language. from video input. Due to the importance of non-manual features (e.g., facial expressions and upper-body movements), many approaches also incorporate these regions to improve recognition accuracy. Some CSLR methods are further extended to support gloss-based sign language translation (SLT) by adding a gloss-to-text module [56, 215, 27].
@@ 6.1 RGB-based
Most state-of-the-art RGB-based Continuous Hand Gesture Recognition (CHGR) models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction [73, 122, 62, 74, 197, 70, 75, 72, 76]. This standard architecture was introduced in [31]. This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries [2].
Across the various methods that extend the initial three-stage pipeline, six broad enhancement strategies emerge: 1) Auxiliary Supervision & Multi‐Task Objectives, which enrich the standard three‐stage pipeline with extra losses and segmentation tasks; 2) Specialized Temporal Modeling, replacing or augmenting the 1D‐CNN with learned pooling, multi-scale fusion, or hierarchical boundary detectors; 3) Attention & Fusion Modules, inserting correlation, self‐emphasis, or transformer‐fusion blocks into the visual encoder; 4) Backbone Replacements & Hybrids, swapping 2D-CNNs for vision transformers or hybrid CNN–attention architectures; 5) Incorporating Linguistic Priors, aligning visual features with gloss embeddings via cross-modal decoders or contrastive losses; and 6) Data Scarcity & Multilingual Strategies, leveraging self-supervision, prototype memories, and automatic dictionary construction to overcome limited annotations and support multiple sign languages.
@@
Many RGB‐based CHGR methods build on the canonical 2DCNN + 1DCNN + BiLSTM architecture by introducing auxiliary losses and multi‐task objectives. For example, VAC‐CSLR [122] augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig. 12). Similarly, SMKD [62] adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR [213] enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
@@
Figure 12: An overview of the VAC-CSLR [122]. The feature extractor provides local information, that is processed by the subsequent alignment module to predict the gloss sequence. $F_{a}$ and $F_{b}$ denote auxiliary classifiers, while $L_{VE}$, $L_{VA}$ and $L_{CTC}$ the VE, VA and CTC loss, respectively. This training setting constitutes the baseline for most of the proposed methods.
@@
To better capture local temporal patterns, several works replace or enhance the standard 1D-CNN. Temporal Lift Pooling (TLP) [73] adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels. mLTSF‐Net [192] couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales. HS‐I3D [190] extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
@@
Other methods inject lightweight attention and correlation blocks into the spatial encoder. CorrNet [74] adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ [76] further replaces its correlation operator with a more efficient variant and inserts an attention layer. SEN‐CSLR [75] introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead. Convolutional Transformer Fusion Blocks (CTFB) [60] fuse intermediate features across 3DCNN branches to enhance cross‐modal integration.
A growing number of approaches swap the 2D-CNN for transformer‐based encoders or hybrid CNN–attention modules. CLIP‐SLA variants [14] leverage a frozen CLIP visual encoder, augmented with LoRA or MLP adapters and an auxiliary classifier for early supervision. Swin‐MSTP [15] pairs a Swin transformer with an MS‐TCN that employs TLP and dilated convolutions, followed by a BiLSTM for global context. TCNet [111] enhances a CNN backbone with trajectory and correlation modules, while SlowFastSign [4] modifies Slow‐Fast‐101 with bidirectional fusion pathways and auxiliary temporal classifiers to refine feature exchange. Most CSLR methods still rely on CNN backbones and thus struggle to model correlations across spatial regions. To address this, SignGraph [^3] Be aware of an IDHGR method with the same name. [47] treats each frame’s patches as nodes in a dynamic graph and alternates between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules to capture intra‐frame and inter‐frame relationships; a multiscale variant, MultiSignGraph, further enhances spatial–temporal context (Fig. 13).
Figure 13: An illustration of the SignGraph [47]. The RGB frames are split into patches. Dynamic graphs are utilized to represent the relations between intra-frame and inter-frame patches, that constitute their nodes.
@@
To exploit gloss co-occurrence and language structure, some models fuse linguistic embeddings with visual features. SignBERT‐HKSL [212] combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and iteratively fine‐tuning with masked outputs. C2ST [197] replaces the CTC decoder with a Conditional Gloss Decoder (CGD) that models gloss dependencies, and aligns visual and gloss embeddings via cross‐modal match loss. GPGN [58] introduces a Cross‐Modal Match Loss between BERT‐derived gloss features and PDC‐TFE–extracted visual embeddings, followed by a BiGRU for sequence modeling (see Fig. 14). CVT‐SLR [205] pretrains an asymmetric VAE on a pseudo gloss‐to‐gloss task, then bridges the VAE and visual encoder with a video–gloss adapter and contrastive alignment loss (see Fig. 15). Fully transformer‐based PiSLTRc [191] applies content‐aware neighborhood gathering and position‐informed attention for end‐to‐end CSLR. Authors in [72] also introduce a prior-aware cross-modality augmentation that trains a transformer to reconstruct masked gloss sequences and then uses it to generate pseudo video–text pairs by guiding gloss substitutions and insertions.
Figure 14: Authors in GPGN [58] leveraged the linguistic prior from the pretrained BERT. The Cross-Modal Match Loss is introduced, in order to maximize the correlation between the gloss and the visual embeddings.
Figure 15: The approach proposed by authors in CVT-SLR [205]. In the first training stage, a VAE is pretrained on the gloss2gloss pseudo-translation task in order to gain textual knowledge. In the second stage, an adopter is placed between the VAE and the visual encoder, while the consistency of the visual and temporal encoder is maintained via contrastive learning.
@@
Finally, data augmentation and cross‐lingual strategies help alleviate limited annotations. Fangyun Wei et al. [188] build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space. CSGC [143] maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.
@@ 6.2 Pose-based
Several works exploit structured keypoint information by dividing the body into semantic regions such as the face, hands, and upper body. MSKA-SLR [56] processes each region independently via attention-enhanced temporal convolutions across four branches. The resulting features are fused to decode gloss sequences using a CTC loss, complemented by frame-level self-distillation. Similarly, CoSign [83] employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency. Their dual-stream variant integrates joint positions and motion features using multi-level fusion to capture both static and dynamic cues.
Temporal localization of gestures has also been explored in the literature. TY-Net [39] draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
Beyond joint coordinates, pose heatmaps serve as an alternative input modality. The multi-cue framework in [210] combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head. The FA-STGCN model [45] further improves robustness to motion scale and frequency variability. It normalizes skeleton inputs and applies alternating spatial and temporal GCNs. Spatial modules learn latent joint relationships, while temporal processing is split into global and local paths—extracting coarse-grained and fine-grained motion patterns, respectively.
Finally, in the context of autonomous driving, Wang et al. [185] developed a two-stage pose-based system to classify road intersection gestures. The first stage extracts geometric features and encodes them via LSTM layers. The second stage uses a dual-stream setup, where one branch processes geometric descriptors and the other uses a ConvLSTM to handle co-occurrence matrices. Features are fused with directional orientation indicators for final classification.
@@ 6.3 Multimodal
Multimodal approaches for CHGR combine multiple input types—such as RGB frames, pose heatmaps, depth maps, and optical flow—to mitigate the limitations of unimodal methods and capture complementary information from different data sources.
@@
A prominent example is TwoStream-SLR [27], a dual-stream architecture designed for both signed language translation (SLT) and CSLR. It processes RGB and pose heatmaps in parallel using the first four blocks of S3D in each branch, with bidirectional lateral connections enabling multi-level information exchange. Additional supervision is provided through Sign Pyramid Networks and auxiliary prediction heads (video, keypoint, and joint). The model is trained using self-distillation, minimizing KL-divergence between the outputs of the three heads. An illustration of the architecture is shown in Fig. 16. This model has also been applied to ISLR [215], using a sliding window mechanism and leveraging the same framework to extract sign boundaries and construct training dictionaries.
@@
Figure 16: Overview of the TwoStream-SLR [27]. Eight S3D blocks constitute the dual-stream multimodal architecture. Bidirectional Lateral Connection is utilized to exchange the information produced by each branch in a multi-level manner. The Sign Pyramid Networks, Video and Keypoint head are actins as auxiliary classifiers, providing extra supervision. Self-distillation is also employed to share the knowledge between the two streams.
@@
Other CSLR models have adapted multimodal ISLR backbones. Uni-Sign [102] can be used directly for CSLR, while SignBERT+ [71] can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.

### v2

@@ 6 Methods for Continuous Dynamic HGR
In contrast to IDHGR, where each sample contains at most one gesture, Continuous Hand Gesture Recognition (CHGR) deals with unsegmented gesture sequences. The goal is to recognize gestures from untrimmed videos and generate the corresponding gesture sequence. CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss (Fangyun_Wei_ICCV; CTC_graves2006connectionist).
While CHGR has been explored in domains like autonomous driving (FA-STGCN; CTCX_WANG2022123), most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss. CSLR aims to produce a sequence of glosses [^1] Glosses are atomic lexical units in sign language, analogous to words in spoken language. from video input. Due to the importance of non-manual features (e.g., facial expressions and upper-body movements), many approaches also incorporate these regions to improve recognition accuracy. Some CSLR methods are further extended to support gloss-based Sign Language Translation (SLT) by adding a gloss-to-text module (MSKA_ELSEVIER; zuo2024onlinecontinuoussignlanguage; NEURIPS2022_6cd3ac24_HEATMAPS).
@@ 6.1 RGB-based
Most state-of-the-art RGB-based CHGR models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction (TemporalLiftPooling; VAC-CSLR; SMKD; CorrNet; C2ST; HezhenHu_TMM; SEN-CSLR; HezhenHu-JunfuPu; corrnetplus_hu2024corrnetsignlanguagerecognition). This standard architecture was introduced in (first_2dcnn_tcn_bilstm_TMM_2019). This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries (ADADOGLOU).
@@
Many RGB‐based CHGR methods build on the canonical 2DCNN + 1DCNN + BiLSTM architecture by introducing auxiliary losses and multi‐task objectives. For example, VAC‐CSLR (VAC-CSLR) augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig. 7). Similarly, SMKD (SMKD) adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR (C2SLR) enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
@@
Figure 7: An overview of the VAC-CSLR (VAC-CSLR). The feature extractor provides local information, that is processed by the subsequent alignment module to predict the gloss sequence. $F_{a}$ and $F_{b}$ denote auxiliary classifiers, while $L_{VE}$, $L_{VA}$ and $L_{CTC}$ the VE, VA and CTC loss, respectively. This training setting constitutes the baseline for most of the proposed methods.
@@
To better capture local temporal patterns, several works replace or enhance the standard 1D-CNN. Temporal Lift Pooling (TLP) (TemporalLiftPooling) adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels. mLTSF‐Net (mLTSF-Net) couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales. HS‐I3D (HS-I3D) extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
@@
Other methods inject lightweight attention and correlation blocks into the spatial encoder. CorrNet (CorrNet) adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ (corrnetplus_hu2024corrnetsignlanguagerecognition) further replaces its correlation operator with a more efficient variant and inserts an attention layer. SEN‐CSLR (SEN-CSLR) introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead. STNet (STNet-wang2025continuous) applied frame-wise loss to the 2DCNN backbone output to enhance feature extraction and inserted multiple Spatial Resonance Modules to facilitate inter-frame modeling.
Many recent approaches replace 2DCNNs with transformer-based or hybrid CNN/attention modules. CLIP-SLA variants (CLIP_SLA_alyami2025clip_CSLR_2025) use a frozen CLIP visual encoder with LoRA or MLP adapters and an auxiliary classifier for early supervision. Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) combines a Swin transformer with an MS-TCN using TLP and dilated convolutions, followed by a BiLSTM for global context.
TCNet (TCNet_Lu_Salah_Poppe_2024) augments a CNN backbone with trajectory and correlation modules, while SlowFastSign (SlowFastSign) modifies Slow-Fast-101 with bidirectional fusion and auxiliary temporal classifiers. To better model spatial correlations, SignGraph (SignGraph_CVPR) treats frame patches as nodes in a dynamic graph, alternating between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules; its multiscale variant, MultiSignGraph, further improves spatiotemporal context.
CORE (CORE-CSLR-zhang2025core) proposed a different approach aiming to leverage the relations between regions within an input frame, deploying an off-the-shelf pose estimator to detect fine-grained and contextual cues; the corresponding frame region are fed into a 2DCNN and the extracted features are fused using a GNN to obtain the final embedding. These per-frame representations pass through a BiLSTM to perform global modeling. SD-Net (SD-Net-CSLR-gao2025structure) enhanced temporal modeling by using three parallel branches. Each branch consists of a two-stage temporal context module that captures both frame-wise and gloss-wise features. Contrastive loss is applied between two of the branches, based on the proposed Entity and Template scheme. Authors in (diffusion-CSLR-geng2026no) formulated CSLR as a cross-modal alignment task using a diffusion model to generate the glosses aiming to mitigate the weak supervision, achieving superior performance. DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) replaced the BiLSTM with a custom contextual model, consisting of memory and feed-forward blocks. Furthermore, the 2DCNN backbone was enhanced by adding a Spatial-Channel Attention (SCA) mechanism, supervised by ground-truth heatmaps, provided by an off-the-shelf pose estimator.
@@
Some models integrate linguistic embeddings with visual features to leverage gloss co-occurrence and language structure. SignBERT-HKSL (SignBERT-HKSL) combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and fine-tuning iteratively on masked outputs. C2ST (C2ST) replaces the CTC decoder with a Conditional Gloss Decoder (CGD) to model gloss dependencies and aligns visual and gloss embeddings via cross-modal match loss. GPGN (GPGN) similarly applies a cross-modal match loss between BERT-derived gloss features and PDC-TFE visual embeddings, followed by a BiGRU for sequence modeling.
CVT-SLR (CVT-SLR) pretrains an asymmetric VAE on a pseudo gloss-to-gloss task, then connects it to a visual encoder with a video–gloss adapter and contrastive alignment loss. Fully transformer-based PiSLTRc (PiSLTRc) uses content-aware neighborhood gathering and position-informed attention for end-to-end CSLR, while (HezhenHu-JunfuPu) introduces prior-aware cross-modality augmentation, training a transformer to reconstruct masked gloss sequences and generate pseudo video–text pairs via guided gloss substitutions and insertions.
@@
Finally, data augmentation and cross‐lingual strategies help alleviate limited annotations. Fangyun Wei et al. (Fangyun_Wei_ICCV) build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space. CSGC (CSGC_Rao_Sun_Wang_Wang_Zhang_2024) maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.
@@ 6.2 Pose-based
Several works exploit structured keypoint information by dividing the body into semantic regions such as the face, hands, and upper body. MSKA-SLR (MSKA_ELSEVIER) processes each region independently via attention-enhanced temporal convolutions across four branches. The resulting features are fused to decode gloss sequences using a CTC loss, complemented by frame-level self-distillation. Similarly, CoSign (CoSign) employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency. Their dual-stream variant integrates joint positions and motion features using multi-level fusion to capture both static and dynamic cues.
Temporal localization of gestures has also been explored in the literature. TY-Net (TY-Net) draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
Beyond joint coordinates, pose heatmaps serve as an alternative input modality. The multi-cue framework in (MULTI-CUE) combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head. The FA-STGCN model (FA-STGCN) further improves robustness to motion scale and frequency variability. It normalizes skeleton inputs and applies alternating spatial and temporal GCNs. Spatial modules learn latent joint relationships, while temporal processing is split into global and local paths—extracting coarse-grained and fine-grained motion patterns, respectively.
Finally, in the context of autonomous driving, Wang et al. (CTCX_WANG2022123) developed a two-stage pose-based system to classify road intersection gestures. The first stage extracts geometric features and encodes them via LSTM layers. The second stage uses a dual-stream setup, where one branch processes geometric descriptors and the other uses a ConvLSTM to handle co-occurrence matrices. Features are fused with directional orientation indicators for final classification.
@@ 6.3 Multimodal
Multimodal CHGR approaches combine inputs such as RGB, pose heatmaps, depth, etc., to overcome unimodal limitations and capture complementary information.
@@
A key example is TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS), a dual-stream architecture for CSLR and SLT. It processes RGB and pose heatmaps in parallel via the first four S3D blocks in each branch, with bidirectional lateral connections for multi-level information exchange. Sign Pyramid Networks and auxiliary heads (video, keypoint, joint) provide additional supervision, while self-distillation minimizes KL-divergence across outputs. Fig. 8 illustrates the architecture. The same framework has been applied to ISLR (zuo2024onlinecontinuoussignlanguage) using a sliding window to extract sign boundaries and construct training dictionaries.
@@
Figure 8: TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) is a dual-stream multimodal architecture built from eight S3D blocks. Bidirectional Lateral Connections enable multi-level information exchange between branches. Auxiliary classifiers, including the Sign Pyramid Network, Video Head, and Keypoint Head, provide additional supervision, while self-distillation transfers knowledge between the two streams.
@@
Other CSLR models have adapted multimodal ISLR backbones. Uni-Sign (UNISIGNli2025) can be used directly for CSLR, while SignBERT+ (SignBERT+) can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.